How do u solve this problem
A top-quality fax-copying machine is purchased for $5800. After t years its value is given by the exponential function V(t)=5800(.8)to the power of t .
The company decides to replace the machine when its value declined to $500. After how long will the machine be replaced?

Answer by stanbon(48535)

A top-quality fax-copying machine is purchased for $5800. After t years its value is given by the exponential function V(t)=5800(.8)to the power of t .
The company decides to replace the machine when its value declined to $500. After how long will the machine be replaced?
-----------------
EQUATION:
500=5800(0.8)^t
Divide by 5800 to get:
5/58 = (0.8)^t
Take the log of both sides:
-1.064457 = tlog(0.8)
-1.064457 = t(-0.09691)
t=10.9839 yrs
===================
Cheers,
Stan H.